# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust in Healthcare

### v2

@@ 1 Introduction
The deployment of AI-driven tools in the healthcare domain has been widespread (see Tjoa and Guan (2021) for an overview), in areas from image classification for brain tumour detection Legastelois et al. (2023) to clinical decision support Domínguez et al. (2023). However, to foster trust in users, the underlying “black-box” AI models need to be supplemented with explanations with consideration of the types of stakeholders the users represent Langer et al. (2021). Explainable AI (XAI) has risen to prominence in recent years as a means of providing users with the explanations they need to be able to trust the outputs of AI models. The ways in which explainability may facilitate trust in users has been studied extensively Ferrario and Loi (2022); Ma et al. (2023), with the general consensus leaning towards it playing an important role in this bond between humans and machines. However, what remains to be seen is which forms of XAI are best suited in individual applications of AI models Hamon et al. (2021). Further, given that the explanatory requirements within healthcare are different from those in general Legastelois et al. (2023), such tailored deployments are especially important in this domain Corti et al. (2024).
@@
One prominent XAI method is feature attribution (FA, see Bodria et al. (2023) for an overview). FA explanations assign each feature for a given input an attribution value denoting its importance towards the output. However, FA methods alone do not account for the human factors involved in enabling trustworthy AI. To evaluate how well XAI methods perform wrt these factors are human-centric metrics such as (perceived) comprehensibility Schmude et al. (2023), i.e. how well users (think they) understand explanations, and trust Jacovi et al. (2021), i.e. how much users trust explanations. However, only a limited number of studies, e.g. Szymanski et al. (2021); Bertrand et al. (2023); Schmude et al. (2023), have considered the effect of the format, rather than the content, of explanations on these metrics.
@@
One recent study Kostopoulou et al. (2022) investigates whether predictions from a cancer risk prediction model, namely the QCancer algorithm Hippisley-Cox and Coupland (2015), have a beneficial effect in supporting the referral decisions of UK-based General Practitioners (GPs). The study found that the inclination of the GPs on whether to refer the patient changed 26% of the time, while their decisions switched entirely only 3% of the time. However, despite the reluctance of GPs to trust and adopt advice from the algorithm, the changed decisions were more consistent with established clinical guidelines, highlighting the potential of this approach to decision support. In a follow-up study, Pálfi et al. Pálfi et al. (2023) found that providing a naive, chart-based FA explanation that visualised the relative weights (based on the regression coefficients of the model) of the features for each prediction did not improve the trust in and uptake of the algorithm, speculating that this could be related to the GPs’ difficulty in comprehending the explanations.
@@
In this paper, we set out to address some of the unanswered questions highlighted by Pálfi et al. Pálfi et al. (2023), broadening from healthcare practitioners to include potential patients also. Namely, we undertook a study involving different stakeholders, aiming to shed light on differing degrees of (perceived) comprehensibility and trust affected in users when they receive explanations of different contents and formats. We do this, in collaboration with healthcare practitioners, by deploying different, well-known FA explanation methods, SHAP Lundberg and Lee (2017) and Occlusion-1 Ancona et al. (2018), for the QCancer algorithm. We undertook two user studies, each with a different group of participants, one sampled from the general population and one from students in medical training, to examine the effect of different stakeholders’ expertise and explanation goals Delaney et al. (2023), which are important factors in the evaluation of explanations Kim et al. (2023); Ehsan et al. (2024). The participants were asked to take part in a survey to assess their (perceived) comprehensibility of and trust in explanations for the outputs of the QCancer algorithm on fictitious patients described by vignettes from Pálfi et al. (2023). Moreover, we created a novel objective measure of comprehension (differentiating from perceived comprehension) to control for any potential reporting biases. Differently from existing studies, we control for the format of explanations (chart or text) to examine their effect on this paradigm. We thus aim to explore the effect of explanation content and format on users’ comprehension and trust via the following research questions (RQs), all considered within the healthcare context:
@@ 2 Related Work
Some works, e.g. Bertrand et al. (2023), also assess FA explanations in text and chart formats, finding that the former increase trust relative to the latter, possibly due to their simplicity. Others have reached similar conclusions, e.g. Szymanski et al. Szymanski et al. (2021) assess visual and textual explanations wrt comprehensibility, controlling for user expertise, as we do, but finding that lay users prefer visual explanations while often misinterpreting them compared to textual ones. Finally, Schmude et al. Schmude et al. (2023) evaluate user comprehension of textual, dialogical and interactive explanations, finding that the former result in worse comprehension and lower preferences from users.
Many works show the importance of expertise with regard to user comprehension and trust, e.g. Wang and Yin (2023), and call for explanations to be designed to suit the needs of individuals Bhatt et al. (2020) and make sense to them Byrne (2023). Other works validate our design choices, e.g. our decision to differentiate between self-reported and objective comprehensibility Cheng et al. (2019). Our approach is a step towards the effective design of multi-modal explanations, e.g. as in Szymanski et al. (2023) in the healthcare domain.
@@
XAI is often considered within healthcare as a tool to enhance trust in and uptake of prediction models Tjoa and Guan (2021). Lebovitz et al. Lebovitz et al. (2022) conducted a user study in which AI was deployed by radiologists to predict breast cancer, lung cancer and bone age: they found that trust in the AI models’ predictions, and thus incorporation into the radiologists’ judgements, were highly dependent on their department, pointing to customised XAI as a potential solution. Moreover, a recent vignette-based user study found promising results Naiseh et al. (2023) in XAI’s deployment in healthcare, finding the AI to be more comprehensible when they received an explanation compared to when they did not. Also, Lee et al. Lee et al. (2024) showed that interactive explanations increased healthcare professionals’ trust in AI, along with their decision accuracy.
Some studies obtained less promising results. For instance, in a user study with healthcare providers, Panigutti et al. Panigutti et al. (2022) assessed the effect of explanations on trust in AI operating a clinical decision support system. They showed that users preferred explanations from the AI model, though the user satisfaction with the explanations themselves was low. Wysocki et al. Wysocki et al. (2023) found that when receiving explanations for COVID-19 risk predictions, users found FA charts easy to interpret but there was no evidence that providing them increased trust in and satisfaction with the AI, compared to when receiving predictions without explanations. A vignette-based study with intensive care doctors found no benefit of an FA explanation over a prediction only option for prescription decisions Nagendran et al. (2023). Finally, a mixed-methods, vignette-based study found that chart-based FA and text-based counterfactual explanations failed to meet the expectations of physicians in reducing uncertainty around heart disease diagnosis Lesley and Hernández (2024).
In summary, few works test XAI’s benefits for healthcare with user studies, with mixed results, which is a known deficiency in the field Hauser et al. (2022); Sadeghi et al. (2024).
@@ 3 Methodology
We adapted twenty fictitious vignettes (including one practice vignette) from Pálfi et al. (2023), describing hypothetical patients presenting at the primary care (i.e. their GP) with symptoms and risk factors suggestive of gastro-oesophageal cancer. We chose to focus on one type of cancer here for simplicity, though the QCancer algorithm and the original vignettes concern multiple types. Each vignette begins with a table of patient demographics and some risk factors, and continues with a textual description of the patient’s symptoms, other risk factors and the results of some relevant examinations. All of the presented factors were predictive of gastro-oesophageal cancer according to the QCancer algorithm. Then, we provided the estimate of the patient’s risk of gastro-oesophageal cancer calculated by the QCancer algorithm with an explanation of this prediction and information on how the explanation should be interpreted. An example of a vignette, with the output of the QCancer algorithm shown for gastro-oesophageal cancer, is given in the following.
@@
To define the explanations that we deliver to users we assume a simple, single-label classification problem comprising a classifier $f:\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}\rightarrow[0,1]$, where $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ represent the domains for each of the $k$ features ($k=15$ in our setting, see Table 1). Then, for any input $\mathbf{x}\in\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}$, $f(\mathbf{x})\in[0,1]$ is the predicted probability of the class given the current values of the input features. This represents the QCancer algorithm, where the features, e.g. age, are inputs to the algorithm and the output is the predicted risk of gastro-oesophageal cancer. We refer to the $i$th value of any input $\mathbf{x}$ as $x_{i}$, where $i\in\{1,\ldots,k\}$. In order to define the FA explanations, we first introduce a simplified version of any given input $\mathbf{x}$ as $\mathbf{x}^{\prime}$, where there exists a mapping function $h_{\mathbf{x}}$ such that $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$, i.e. returning the original input given the simplified version. We then let, as in Lundberg and Lee (2017), $\mathbf{x}^{\prime}\in\{0,1\}^{M}$, where $M$ is the number of simplified input features (note that $M=k$ throughout this paper). Then, if $x_{i}^{\prime}=1$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the current value of the feature, while if $x_{i}^{\prime}=0$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the feature’s baseline value, i.e. a pre-defined default value. We set the baseline values (referred to as baseline patients in the explanations) to one of two sets of values representing healthy males and females depending on the patient’s gender, as indicated in Table 1, which were used for the studies.
One of the most prominent of all FA methods is SHAP Lundberg and Lee (2017), where each feature of an input to a model is assigned an importance value computed from a game-theoretical approximation of its contribution.
@@
We then introduce chart-based SHAP (SC) explanations, which are the explanations computed by SHAP expressed as a percentage in chart format, as exemplified in Figure 1 (bottom). [^3] We will share the link to our visualisation tool on acceptance. For all explanations, the displayed features were the six with the highest magnitudes. Note that all explanations were accompanied by a textual description of the explanation, as shown in Figure 1 (top). In the example explanation in Figure 1, we can see that the top five features increased the risk of gastro-oesophageal cancer, with difficulty swallowing and its value of yes being the most important. The last feature, BMI, with value 30.643, reduced the risk but had a smaller magnitude, and thus importance, as calculated by SHAP.
@@
We then selected the Occlusion-1 method Ancona et al. (2018) for its simplicity in hopes of achieving more comprehensibility and trust in users. This method takes the set of baseline feature values (as in SHAP explanations) and assigns any feature an attribution corresponding to the difference in the prediction if the feature’s value is changed to the baseline value (with all other features’ values remaining unchanged).
@@
Note that, while these explanations do not explicitly mention datapoints, each feature’s attribution in explanations implicitly refers to a modified datapoint with that feature set to the baseline value, with all other features remaining unchanged. Thus, while this differs from counterfactual and semi-factual explanations in the literature (see Guidotti (2022); Aryal and Keane (2023), respectively, for recent overviews), which are based on single modified datapoints, it still gives an actionable representation of (six) modifications users could make to the input to achieve a different output (as in Albini et al. Albini et al. (2020)). We thus tentatively predicted that this inherent counterfactual nature of Occlusion-1 explanations may give more amenability to humans, as observed for counterfactual explanations by Miller Miller (2019), which strengthened our prediction of the rejection of the null hypotheses for RQ1-3.
@@
We assessed the RQs in two different samples: one from the general population where a limited number of participants had medical experience, representing stakeholders in the form of patients (Study 1), and an expert sample where participants had some medical training and experience, representing stakeholders in the form of healthcare practitioners (Study 2). We created two Qualtrics [^5] https://www.qualtrics.com surveys using the following structure (see the Supplementary Material for an example):
@@
For Study 1, we aimed to recruit 100 participants from the general population, who were over 18 and spoke English via the online research platform Prolific [^6] Prolific (https://www.prolific.com) was chosen for its superior data quality when compared with other platforms Peer et al. (2022).. Upon completion, the participants were paid an average of \$4.50 for their time based on the platform’s standard reimbursement rates of \$21.20 per hour. Data collection was anonymous during Study 1. In Study 2, we aimed to recruit at least 50 students in medicine from UK-based universities. The potential participants were approached via student mailing lists, and were invited via a link to the survey in an email, which also contained the participant information sheet with details about the study. All students were at least 18 years of age and English-speaking as they were attending university in the UK. At the end of Study 2, participants were asked to provide their university email address, to which we sent a £10 voucher for the online retailer Amazon to reimburse them. [^7] The anonymised and pseudononymised data will be shared via Zenodo (https://zenodo.org) on acceptance.
@@
We ran the statistical analyses in R (version 4.3.1). We built generalised linear and linear mixed-effects regression models with the lme4 R package Bates et al. (2015) to analyse our categorical and interval variables, respectively. For the generalised linear regression models, we only included random intercepts by participants, and for the linear regression models, we included random intercepts by participants and vignettes. For the generalised linear and the linear regression models we report Odds Ratios (ORs) and raw regression slopes (bs), respectively. We also report 95% CIs for each regression coefficient. To test our hypotheses, we used a p-value of 0.05.
We ran two types of analyses for RQ2 and RQ3. First, we compared the SHAP to a single Occlusion-1 explanation group, merging OC and OT explanations, for a general test of explanation content. Second, we made pairwise comparisons of explanations to examine the effects of explanation content, format and their interaction. We did not conduct formal analyses of the free-text responses of general comments and the justifications about trust towards the explanations due to the low response rates (15% across questions and both studies).

### v3

@@ 1 Introduction
The deployment of AI-driven tools in the healthcare domain has been widespread (see [42] for an overview), in areas from image classification for brain tumour detection [26] to clinical decision support [12]. However, to foster trust in users, the underlying “black-box” AI models need to be supplemented with explanations with consideration of the types of stakeholders the users represent [23]. Explainable AI (XAI) has risen to prominence in recent years as a means of providing users with the explanations they need to be able to trust the outputs of AI models. The ways in which explainability may facilitate trust in users has been studied extensively [15, 29], with the general consensus leaning towards it playing an important role in this bond between humans and machines. However, what remains to be seen is which forms of XAI are best suited in individual applications of AI models [17]. Further, given that the explanatory requirements within healthcare are different from those in general [26], such tailored deployments are especially important in this domain [10].
@@
One of the most prominent types of XAI method is feature attribution (FA, see [7] for an overview). FA explanations assign each feature for a given input an attribution value denoting its importance towards the output. However, FA methods alone do not account for the human factors involved in enabling trustworthy AI. To evaluate how well XAI methods perform wrt these factors are human-centric metrics such as (perceived) comprehensibility [39], i.e. how well users (think they) understand explanations, and trust [19], i.e. how much users trust explanations. However, only a limited number of studies, e.g. [40, 5, 39], have considered the effect of the format, rather than the content, of explanations on these metrics.
@@
One recent study [21] investigates whether predictions from a cancer risk prediction model, namely the QCancer algorithm [18], have a beneficial effect in supporting the referral decisions of UK-based General Practitioners (GPs). The study found that the inclination of the GPs on whether to refer the patient changed 26% of the time, while their decisions switched entirely only 3% of the time. However, despite the reluctance of GPs to trust and adopt advice from the algorithm, the changed decisions were more consistent with established clinical guidelines, highlighting the potential of this approach to decision support. In a follow-up study, Pálfi et al. [34] found that providing a naive, chart-based FA explanation that visualised the relative weights (based on the regression coefficients of the model) of the features for each prediction did not improve the trust in and uptake of the algorithm, speculating that this could be related to the GPs’ difficulty in comprehending the explanations.
@@
In this paper, we set out to address some of the unanswered questions highlighted by Pálfi et al. [34], broadening from healthcare practitioners to include potential patients also. Namely, we undertook a study involving different stakeholders, aiming to shed light on differing degrees of (perceived) comprehensibility and trust affected in users when they receive explanations of different contents and formats. We restrict our focus to FA explanations as in [34], and leave the study of whether our findings translate to other XAI methods to future work. We undertake this study, in collaboration with healthcare practitioners, by deploying different, well-known FA explanation methods, SHAP [28] and Occlusion-1 [2] (two of the most cited works in this area), for the QCancer algorithm in an online system, providing a useful tool for future research. We undertook two user studies, each with a different group of participants, one sampled from the general population and one from students in medical training, to examine the effect of different stakeholders’ expertise and explanation goals [11], which are important factors in the evaluation of explanations [20, 14]. The participants were asked to take part in a survey to assess their (perceived) comprehensibility of and trust in explanations for the outputs of the QCancer algorithm on fictitious patients described by vignettes from [34]. Moreover, we created a novel objective measure of comprehension (differentiating from perceived comprehension) to control for any potential reporting biases. Differently from existing studies, we control for the format of explanations (chart or text) to examine their effect on this paradigm. We thus aim to explore the effect of explanation content and format on users’ comprehension and trust via the following research questions (RQs), all considered within the healthcare context:
@@ 2 Related Work
Some works, e.g. [5], also assess FA explanations in text and chart formats, finding that the former increase trust relative to the latter, possibly due to their simplicity. Others have reached similar conclusions, e.g. Szymanski et al. [40] assess visual and textual explanations wrt comprehensibility, controlling for user expertise, as we do, but find that lay users prefer visual explanations while often misinterpreting them compared to textual ones. Finally, Schmude et al. [39] evaluate user comprehension of textual, dialogical and interactive explanations, finding that the former result in worse comprehension and lower preferences from users.
Many works show the importance of expertise with regard to user comprehension and trust, e.g. [43], and call for explanations to be designed to suit the needs of individuals [6] and make sense to them [8]. Other works validate our design choices, e.g. our decision to differentiate between self-reported and objective comprehensibility [9]. Our approach is a step towards the effective design of multi-modal explanations, e.g. as in [41] in the healthcare domain.
@@
XAI is often considered within healthcare as a tool to enhance trust in and uptake of prediction models [42]. Lebovitz et al. [24] conducted a user study in which AI was deployed by radiologists to predict breast cancer, lung cancer and bone age: they found that trust in the AI models’ predictions, and thus incorporation into the radiologists’ judgements, were highly dependent on their department, pointing to customised XAI as a potential solution. Moreover, a recent vignette-based user study found promising results [32] in XAI’s deployment in healthcare, finding the AI to be more comprehensible when they received an explanation compared to when they did not. Also, Lee et al. [25] showed that interactive explanations increased healthcare professionals’ trust in AI, along with their decision accuracy.
Some studies obtained less promising results. For instance, in a user study with healthcare providers, Panigutti et al. [35] assessed the effect of explanations on trust in AI operating a clinical decision support system. They showed that users preferred explanations from the AI model, though the user satisfaction with the explanations themselves was low. Wysocki et al. [44] found that when receiving explanations for COVID-19 risk predictions, users found FA charts easy to interpret but there was no evidence that providing them increased trust in and satisfaction with the AI, compared to when receiving predictions without explanations. A vignette-based study with intensive care doctors found no benefit of an FA explanation over a prediction only option for prescription decisions [31]. Finally, a mixed-methods, vignette-based study found that chart-based FA and text-based counterfactual explanations failed to meet the expectations of physicians in reducing uncertainty around heart disease diagnosis [27].
In summary, few works test XAI’s benefits for healthcare with user studies, with mixed results, which is a known deficiency in the field [38].
@@ 3 Methodology
We adapted twenty fictitious vignettes (including one practice vignette) from [34], describing hypothetical patients presenting at the primary care (i.e. their GP) with symptoms and risk factors suggestive of gastro-oesophageal cancer. We chose to focus on one type of cancer here for simplicity, though the QCancer algorithm and the original vignettes concern multiple types. Each vignette begins with a table of patient demographics and some risk factors, and continues with a textual description of the patient’s symptoms, other risk factors and the results of some relevant examinations. All of the presented factors were predictive of gastro-oesophageal cancer according to the QCancer algorithm. Then, we provided the estimate of the patient’s risk of gastro-oesophageal cancer calculated by the QCancer algorithm with an explanation of this prediction and information on how the explanation should be interpreted. An example of a vignette, with the output of the QCancer algorithm shown for gastro-oesophageal cancer, is given in the following.
@@
To define the explanations that we deliver to users we assume a simple, single-label classification problem comprising a classifier $f:\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}\rightarrow[0,1]$, where $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ represent the domains for each of the $k$ features ($k=15$ in our setting, see Table 1). Then, for any input $\mathbf{x}\in\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}$, $f(\mathbf{x})\in[0,1]$ is the predicted probability of the class given the current values of the input features. This represents the QCancer algorithm, where the features, e.g. age, are inputs to the algorithm and the output is the predicted risk of gastro-oesophageal cancer. We refer to the $i$th value of any input $\mathbf{x}$ as $x_{i}$, where $i\in\{1,\ldots,k\}$. In order to define the FA explanations, we first introduce a simplified version of any given input $\mathbf{x}$ as $\mathbf{x}^{\prime}$, where there exists a mapping function $h_{\mathbf{x}}$ such that $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$, i.e. returning the original input given the simplified version. We then let, as in [28], $\mathbf{x}^{\prime}\in\{0,1\}^{M}$, where $M$ is the number of simplified input features (note that $M=k$ throughout this paper). Then, if $x_{i}^{\prime}=1$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the current value of the feature, while if $x_{i}^{\prime}=0$, then $\mathbf{x}=h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the feature’s baseline value, i.e. a pre-defined default value. We set the baseline values (referred to as baseline patients in the explanations) to one of two sets of values representing healthy males and females depending on the patient’s gender, as indicated in Table 1, which were used for the studies.
One of the most prominent of all FA methods is SHAP [28], where each feature of an input to a model is assigned an importance value computed from a game-theoretical approximation of its contribution.
@@
We then introduce chart-based SHAP (SC) explanations, which are the explanations computed by SHAP expressed as a percentage in chart format, as exemplified in Figure 1 (bottom).For all explanations, the displayed features were the six with the highest magnitudes. Note that all explanations were accompanied by a textual description of the explanation, as shown in Figure 1 (top). In the example explanation in Figure 1, we can see that the top five features increased the risk of gastro-oesophageal cancer, with difficulty swallowing and its value of yes being the most important. The last feature, BMI, with value 30.643, reduced the risk but had a smaller magnitude, and thus importance, as calculated by SHAP.
@@
We then selected the Occlusion-1 method [2] for its simplicity in hopes of achieving more comprehensibility and trust in users. This method takes the set of baseline feature values (as in SHAP explanations) and assigns any feature an attribution corresponding to the difference in the prediction if the feature’s value is changed to the baseline value (with all other features’ values remaining unchanged).
@@
Note that, while these explanations do not explicitly mention datapoints, each feature’s attribution in explanations implicitly refers to a modified datapoint with that feature set to the baseline value, with all other features remaining unchanged. Thus, while this differs from counterfactual and semi-factual explanations in the literature (see [16, 3], respectively, for recent overviews), which are based on single modified datapoints, it still gives an actionable representation of (six) modifications users could make to the input to achieve a different output (as in Albini et al. [1]). We thus tentatively predicted that this inherent counterfactual nature of Occlusion-1 explanations may give more amenability to humans, as observed for counterfactual explanations by Miller [30], which strengthened our prediction of the rejection of the null hypotheses for RQ1-3.
@@
All explanations in Figures 1 to 3 were visualised with the purpose-built online tool [^5] https://qcancer-explanation.herokuapp.com/# that we developed for explaining the prdictions of the QCancer algorithm, the user interface for which is shown in Figure 4. SHAP explanations remain optional in the system due to their additional computation time. It should be noted that more features are included here (compared with Table 1) as the online tool can be used to explain QCancer predictions for all supported types of cancer, not only gastro-oesophageal cancer as is our focus in this paper. Our hope is for this tool to help support future (user) studies of XAI for the QCancer algorithm.
Figure 4: User interface for our online tool for visualising different forms of explanation for the QCancer algorithm, showing the OT explanation for the patient from Example 1 and the next highest predicted type of cancer (lung) after gastro-oesophageal.
@@
We assessed the RQs in two different samples: one from the general population where a limited number of participants had medical experience, representing stakeholders in the form of patients (Study 1), and an expert sample where participants had some medical training and experience, representing stakeholders in the form of healthcare practitioners (Study 2). We created two Qualtrics [^6] https://www.qualtrics.com surveys using the following structure (see the Supplementary Material [^7] The Supplementary Material is available at https://www.dropbox.com/scl/fi/b2vnsdisbzqozz3megbi5/sm.pdf?rlkey=fst
dro2smseq40iakcsmzqv6e&dl=0
for an example):
@@
For Study 1, we aimed to recruit 100 participants from the general population, who were over 18 and spoke English via the online research platform Prolific [^8] Prolific (https://www.prolific.com) was chosen for its superior data quality when compared with other platforms [36].. Upon completion, the participants were paid an average of \$4.50 for their time based on the platform’s standard reimbursement rates of \$21.20 per hour. Data collection was anonymous during Study 1. In Study 2, we aimed to recruit at least 50 students in medicine from UK-based universities. The potential participants were approached via student mailing lists, and were invited via a link to the survey in an email, which also contained the participant information sheet with details about the study. All students were at least 18 years of age and English-speaking as they were attending university in the UK. At the end of Study 2, participants were asked to provide their university email address, to which we sent a £10 voucher for the online retailer Amazon to reimburse them. [^9] The anonymised and pseudononymised data will be shared via Zenodo (https://zenodo.org) on acceptance.
Participation in both studies was voluntary. Participants were asked to provide informed consent after reading an information sheet and they could withdraw at any time prior to completion, simply by closing the survey window, with their data deleted permanently. We informed the participants that once they had completed the survey, they could no longer withdraw and we would keep their data. All the participant responses were anonymised once payment was made (i.e. we deleted their e-mail addresses from the dataset). We received ethics approval from the institution leading these studies.
@@
We ran the statistical analyses in R (version 4.3.1). We built generalised linear and linear mixed-effects regression models with the lme4 R package [4] to analyse our categorical and interval variables, respectively. For the generalised linear regression models, we only included random intercepts by participants, and for the linear regression models, we included random intercepts by participants and vignettes. For the generalised linear and the linear regression models we report Odds Ratios (ORs) and raw regression slopes (bs), respectively. We also report 95% CIs for each regression coefficient. To test our hypotheses, we used a p-value of 0.05.
We ran two types of analyses for RQ2 and RQ3. First, we compared SHAP explanations to a single Occlusion-1 explanation group, merging OC and OT explanations, for a general test of explanation content. Second, we made pairwise comparisons of explanations to examine the effects of explanation content, format and their interaction. We did not conduct formal analyses of the free-text responses of the comments and justifications on trust towards the explanations due to the low response rates (15% across questions and both studies).